### v1

@@ 3.4 Enhancing Memory in SSL
In SSL, we mainly discuss Contrastive Learning (CL). CL learns representations of data by focusing on positive samples that are close to each other (such as augmentations from the same image) and excluding negative samples (such as augmentations from different images) (Wu et al., 2024). By promoting the closeness of positive examples and maximizing the separation between negative examples in the latent space, it learns the representation of the data (Gui et al., 2024).
In CL, a relatively simple and practical framework is SimCLR (Chen et al., 2020), which is an easy-to-implement visual representation contrastive learning framework. We mainly discuss the training process of the basic encoder $f(\bullet):\mathbb{R}^{L\times 1}\longrightarrow\mathbb{R}^{H\times 1}$, where $f(\bullet)$ typically uses a ResNet architecture (He et al., 2016). For a sample $x\in\mathbb{R}^{L\times 1}$, two independent data augmentation operators can be sampled to obtain $x_{i},x_{j}\in\mathbb{R}^{L\times 1}$. The basic encoder maps the two data to intermediate representations $f(x_{i}),f(x_{j})$. During the training phase, a projection head $g(\bullet)$ is often used to obtain the outputs $g(f(x_{i})),g(f(x_{j}))$. And training is conducted using the NT-Xent loss to decrease.
@@
Previous work (Feldman, 2020) has explored model memory, where in SL, they compared the model’s probability of a sample before and after the sample’s participation in training. Since in CL, the model primarily learns the intrinsic features of the data without involving data labels in training, we extend this concept to a more general form:
$mem\left(x\right)=\begin{matrix}loss\left(x_{i},x_{j}\right)\\ A\left(D-x\right)\\ \end{matrix}-\begin{matrix}loss\left(x_{i},x_{j}\right)\\ A\left(D\right)\\ \end{matrix}$ (11)
@@
It is worth noting that in CL, if the model fits the sample $x$ well, it implies that the model will judge the two data generated from $x$, $x_{i}$ and $x_{j}$, as positive samples, and the NT-Xent loss function constructed by $x_{i}$ and $x_{j}$ will also decrease. In other words, when the model’s memory of the sample $x$ is enhanced, it means that for $x_{i}$ and $x_{j}$, the model’s output tends to be the same, which is determined by the SimCLR framework. Chen et al. (2020) points out that whether it is a linear or nonlinear projection head, $f(x_{i})$ and $f(x_{j})$ form and maintain more information, and the hidden layers before projection are often richer representations. Therefore, we can explicitly consider that when the sample $x$ is remembered, the basic encoder’s output for $x_{i}$ and $x_{j}$ will be as similar as possible, so the $loss()$ can be represented as:
@@
However, for the first term of Equation (11), we found that in datasets with a sufficiently large number of samples, for most samples, this term is almost uniform. As shown in Figure (4) and (5), we conducted two experiments: a) We randomly selected 50 samples $\{x^{(1)},\ldots,x^{(i)},\ldots,x^{(50)}\}$ and examined the $\begin{matrix}loss(x_{i},x_{j})\\ A(D-B)\\ \end{matrix}$ for each sample. b) To explore the impact of the number of sampled data on the original dataset, we randomly sampled a certain proportion of samples from the entire dataset, denoted as $B_{0}=r\times D_{0}$, where $B$ is the set of samples, and we investigated the relationship between the value of $\mathbb{E}_{x\in B}\begin{matrix}loss(x_{i},x_{j})\\ A(D-B)\\ \end{matrix}$ and the proportionality factor $r$.
@@
According to the experiments, when a single sample is removed, the variation of $\text{loss}(x_{i},x_{j})$ with $A(D-x)$ is not significant. We attribute this to the model having already learned sufficient knowledge from most of the data, and thus is not sensitive to these individual samples. Therefore, we can fully consider that for different data, the variation of $\begin{matrix}loss(x_{i},x_{j})\\ A(D-B)\\ \end{matrix}$ can be neglected. Moreover, as the sampling ratio $r$ increases, this value of $\begin{matrix}loss(x_{i},x_{j})\\ A(D-B)\\ \end{matrix}$ becomes more pronounced, which makes the first term in Equation (11) non-negligible. In extreme cases, the performance of the data pruning method will drop sharply, reflecting the difficulty of data pruning in CL at extreme pruning rates, which will be our future work target.
@@
Where $NX(\bullet)$ represents the NT-Xent loss, and $\beta$ is an adjustable hyperparameter.
@@
0:  Dataset $D$, Epoch $T$, Pruning Ratio $s$, Initial Model $\theta$, Retained Dataset $\hat{D}_{0}=D$.
1:  for $t$ from 0 to $T-1$ do
2:     if $t!=0$ then
3:        Select the retained dataset $\hat{D}_{t}$ based on the pruning rate $s$ and sample scores.
4:     else
5:        $\hat{D}_{t}=D$
6:     end if
7:     Get train sequence $\hat{D}_{t}=\{B_{0},B_{1},,,B_{b_{t}-1}\}$
8:     for $j$ from 0 to $b_{t}-1$ do
9:        Get NX-loss $NX\left(g\left(f\left(B_{j}\right)\right),g\left(f\left(B_{j}\right)\right)\right)$
10:        Update score with Equation (13)
11:        Update model with optimizer
12:     end for
@@ 4 EXPERIMENT
In the following sections, we validate the effectiveness of our theoretical results and the proposed dataset pruning method through experiments. In Section 4.1, we compare EMP with several other baseline methods on SL, including image classification tasks and a range of natural language tasks. In Section 4.2, we verify the high efficiency of EMP on SSL. In Section , we conduct a series of ablation experiments to validate the effectiveness of our theoretical results. In Section 4.4, we analyze the generalization performance of EMP using a one-dimensional linear interpolation method. It is worth noting that other baseline methods include static pruning: CD (Agarwal et al., 2020), Herding (Welling, 2009), DeepFool (Ducoffe and Precioso, 2018), Last Confidence (Coleman et al., ), Glister (Killamsetty et al., 2021), and dynamic pruning: InfoBatch (Qin et al., ), $\epsilon$-greedy, and UCB (Raju et al., 2021). It is important to note that EMP is also a dynamic pruning method, so we mainly compare it with other dynamic pruning methods. To eliminate the influence of randomness, each of our experiments was run 5 times, and the average was taken. Our dataset introduction and experimental details are in the Appendix B and C.
@@ 4.1 Performance of EMP in SL
CIFAR. We conducted comparisons on both CIFAR-10 and CIFAR-100. In this work, similar to InfoBatch (Qin et al., ), we employed an annealing technique, specifically, using the entire dataset for training towards the end of the training process. More details are shown in Section 4.3. We presented results from other methods in recent years in Table (1). EMP achieved lossless performance on CIFAR-100, even slightly higher than the baseline. Notably, EMP significantly surpassed previous static pruning methods and led other dynamic pruning methods at high pruning rates. Specifically, on the CIFAR-100 dataset, EMP exceeded other methods by at least 1.3% under any pruning rate, and even by at least 2.1% at a 70% pruning rate.
@@
CD(Agarwal et al. (2020)) 95.0 94.3 90.8 74.2 72.3 70.3
Herding(Welling (2009)) 92.2 88.0 80.1 73.1 71.8 69.6
K-Center(Sener and Savarese (2018)) 94.7 93.9 90.9 74.1 72.2 70.2
Least Confidence(Coleman et al. ) 95.0 94.5 90.3 74.2 72.3 69.8
Margin(Coleman et al. ) 94.9 94.3 90.9 74.0 72.2 70.2
GraNd-4(Paul et al. (2021)) 95.3 94.6 91.2 74.6 71.4 68.8
DeepFool(Ducoffe and Precioso (2018)) 95.1 94.1 90.0 74.2 73.2 69.8
Craig(Mirzasoleiman et al. (2020)) 94.8 93.3 88.4 74.4 71.9 69.7
Glister(Killamsetty et al. (2021)) 95.2 94.0 90.9 74.6 73.2 70.4
EL2N-2(Toneva et al. ) 94.4 93.2 89.8 74.1 71.0 68.5
EL2N-20(Toneva et al. ) 95.3 95.1 91.9 77.2 72.1 -
DP(Yang et al. ) 94.9 93.8 90.8 77.2 73.1 -
@@
$\epsilon$-greedy(Raju et al. (2021)) 95.2 94.9 94.1 76.4 74.8 75.1
UCB(Raju et al. (2021)) 95.3 94.7 93.9 77.3 75.3 74.8
InfoBatch(Qin et al. ) 95.6 95.1 94.7 78.2 78.1 76.5
EMP(ours) 95.37±$\pm 0.11$ 95.23±$\pm 0.15$ 95.04±$\pm 0.08$ 79.53±$\pm 0.21$ 79.44±$\pm 0.18$ 78.63±$\pm 0.42$
Baseline 95.6±$\pm 0.1$ 79.8±$\pm 0.04$
@@
ImageNet-1K. To explore the performance of EMP on a large-scale dataset, we trained ResNet50 on ImageNet-1K, with results as shown in Table 2. Notably, the UCB and Greedy methods demonstrated performance similar to that of dynamic random pruning, and at high pruning rates, dynamic random pruning surpassed them. We believe that on ImageNet-1K, due to some noisy samples (Beyer et al., 2020), they are difficult to fit correctly, hence LFL tends to select these noisy samples, leading to suboptimal performance of LFL. As expected, on the challenging ImageNet-1K dataset with high fitting difficulty, EMP showed efficient performance both at low and high pruning rates.
@@
EMP(ours) 74.79$\pm$0.15 72.77$\pm$0.11
Baseline 75.36$\pm$0.15
GLUE. In addition to image classification tasks, natural language understanding tasks are also within the scope of our investigation. Specifically, we used the BERT-base pre-trained model (Devlin, 2018) for fine-tuning on the GLUE benchmark (Wang, 2018). The experimental results are reported in Table (3), and it is evident that EMP leads in most cases, even outperforming the case without pruning. Notably, at pruning rates of 50% and 70%, for datasets with a larger number of samples such as MNLI (393K) and QQP (363K), EMP demonstrates good performance. Specifically, at a 70% pruning rate, EMP leads other methods by at least 1% on the MNLI dataset. On the dataset QNLI with a moderate number of samples (108K), EMP also shows superior performance, exceeding other methods by 0.6%-1.9% at a 70% pruning rate. This proves that our method is also effective on natural language tasks.
@@ 4.2 Performance of EMP in SSL
Recently, Large Language Models (LLMs) have demonstrated significant performance (Touvron et al., 2023; Achiam et al., 2023), pretraining on vast datasets to obtain models with general knowledge. Visually, various visual pretrained models (Dehghani et al., 2023; Kim et al., 2021) are also gaining popularity. To explore the performance of EMP in model pretraining, we experimentally verified it on CL using the SimCLR framework (Chen et al., 2020). In this work, we specifically pruned the data during the pretraining phase to investigate the practical effectiveness of EMP.
@@
$\epsilon$-greedyCover (1999) 82.01 79.98 76.91 53.11 49.89 42.72
UCBCover (1999) 81.37 79.24 71.85 52.50 49.34 42.15
InfoBatchRaju et al. (2021) 83.09 81.39 79.05 55.21 50.56 46.01
EMP(ours) 83.19±$\pm 0.33$ 82.16±$\pm 0.37$ 80.59±$\pm 0.28$ 56.37±$\pm 0.41$ 51.99±$\pm 0.52$ 48.21±$\pm 0.47$
Baseline 84.67±$\pm 0.17$ 56.92±$\pm 0.44$
@@ 4.3 Ablation Experiment
It can be seen that the degree of annealing increases with the increase of $\alpha$, but this also brings corresponding additional overhead. We believe that when the model has acquired sufficient knowledge in the non-annealing phase, the positive effects brought by the annealing phase will be less significant. In other words, the annealing phase not only exists as a technique to improve accuracy but also serves as an effective measure of the validity of the data pruning algorithm. Similarly, InfoBatch (Qin et al., ) also uses this technique, and we mainly compare it with this method. Our results are reported in Figure (8). It is not difficult to see that at lower pruning rates, both methods are robust to annealing, and when the pruning rate increases, the shortcomings of LFL become apparent. Although EMP’s accuracy drops without the annealing phase, InfoBatch’s drop is more dramatic. Therefore, we can conclude that under any circumstances, EMP is robust to the annealing technique, which means that EMP enables the model to remember more samples and gain more knowledge during the early to mid-training phase.
@@
Label Noise. Label noise is a common challenge in real-world applications. Enhancing the robustness of algorithms to label noise during the data pruning process is a key issue (Tan et al., 2024). In this section, we investigate the robustness of EMP to label noise by conducting comparative experiments on CIFAR10 and CIFAR100 with synthetic label noise. Specifically, we introduce label noise (Patrini et al., 2017) into the two datasets by randomly replacing the labels of a certain percentage of training data with all possible labels, which mimics the real-life scenario where researchers may misjudge images (Rolnick et al., 2017). Our results are reported in Figure (9). It can be observed that EMP still maintains high performance even with the addition of 20% noise and leads other methods. Notably, the Greedy and UCB methods are not robust to noise, and are even outperformed by random dynamic pruning. We believe that these noisy data are difficult to fit during training, hence their scores remain high, leading to the retention of noisy data and resulting in low model performance. Overall, EMP maintains high performance in the presence of noise, indicating that EMP indeed retains key or general samples, thus demonstrating robustness to noise.
@@ 4.4 Generalization Analysis
To further investigate the effectiveness of EMP, we analyze its generalization capability. We use one-dimensional linear interpolation to examine how EMP affects the loss landscape while enhancing the model’s memory ability. Previous work (Hochreiter and Schmidhuber, 1994, 1997) suggests that better model generalization is associated with flat minima in the loss landscape. Following the method proposed in (Goodfellow et al., 2014), we inspect the loss landscape. Specifically, we assess the performance of models with parameters $(1-\epsilon)\theta_{\text{init}}+\epsilon\theta_{T}$, where $\theta_{\text{init}}$ is the initial model, and $\theta_{T}$ is the converged model after optimization with different dataset pruning algorithms.

### v2

@@ 3.4 Enhancing Memory in SSL
In SSL, we mainly discuss Contrastive Learning (CL). CL learns representations of data by focusing on positive samples that are close to each other (such as augmentations from the same image) and excluding negative samples (such as augmentations from different images) [ref13]. By promoting the closeness of positive examples and maximizing the separation between negative examples in the latent space, it learns the representation of the data [ref14].
In CL, a relatively simple and practical framework is SimCLR [ref15], which is an easy-to-implement visual representation contrastive learning framework. We mainly discuss the training process of the basic encoder $f(\cdot):\mathbb{R}^{L\times 1}\longrightarrow\mathbb{R}^{H\times 1}$, where $f(\cdot)$ typically uses a ResNet architecture [ref16]. For a sample $x\in\mathbb{R}^{L\times 1}$, two independent data augmentation operators can be sampled to obtain $x_{i},x_{j}\in\mathbb{R}^{L\times 1}$. The basic encoder maps the two data to intermediate representations $f(x_{i}),f(x_{j})$. During the training phase, a projection head $g(\cdot)$ is often used to obtain the outputs $g(f(x_{i})),g(f(x_{j}))$. And training is conducted using the NT-Xent loss to decrease.
@@
Previous work [ref11] has explored model memory, where in SL, they compared the model’s probability of a sample before and after the sample’s participation in training. Since in CL, the model primarily learns the intrinsic features of the data without involving data labels in training, we extend this concept to a more general form:
[…]
@@
It is worth noting that in CL, if the model fits the sample $x$ well, it implies that the model will judge the two data generated from $x$, $x_{i}$ and $x_{j}$, as positive samples, and the NT-Xent loss function constructed by $x_{i}$ and $x_{j}$ will also decrease. In other words, when the model’s memory of the sample $x$ is enhanced, it means that for $x_{i}$ and $x_{j}$, the model’s output tends to be the same, which is determined by the SimCLR framework. [ref15] points out that whether it is a linear or nonlinear projection head, $f(x_{i})$ and $f(x_{j})$ form and maintain more information, and the hidden layers before projection are often richer representations. Therefore, we can explicitly consider that when the sample $x$ is remembered, the basic encoder’s output for $x_{i}$ and $x_{j}$ will be as similar as possible, so the $loss()$ can be represented as:
@@
[…]
@@
[…]
@@
Where $NX(\cdot)$ represents the NT-Xent loss, and $\beta$ is an adjustable hyperparameter.
@@
1:Dataset $D$, Epoch $T$, Pruning Ratio $s$, Initial Model $\theta$, Retained Dataset $\hat{D}_{0}=D$.
2:for $t$ from 0 to T1T-1 do
3:  if $t!=0$ then
4:   Select the retained dataset $\hat{D}_{t}$ based on the pruning rate $s$ and sample scores.
5:  else
6:   $\hat{D}_{t}=D$
7:  end if
8:  Get train sequence $\hat{D}_{t}=\{B_{0},B_{1},,,B_{b_{t}-1}\}$
9:  for $j$ from 0 to $b_{t}-1$ do
10:   Get NX-loss $NX\left(g\left(f\left(B_{j}\right)\right),g\left(f\left(B_{j}\right)\right)\right)$
11:   Update score with Equation (13)
12:   Update model with optimizer
@@
14:end for
The limitation of static pruning lies in the fact that methods like Yang et al.[yangdataset] rely on proxy models. This introduces architectural bias - the agent’s "perception" of the importance of data may not be transferred to the target model, thereby undermining generalization.
EMP continuously scores and trims the data at each training checkpoint (Algorithms 1 and 2). By evaluating samples based on the current model state, it retains the data most relevant to the model’s instantaneous learning stage. This ensures that the training subset evolves along with the model and maintains consistency with its generalization trajectory.
The dynamic approach of EMP transforms pruning from a one-time filtering to a continuous feedback loop between the model state and data utility. By aligning subset selection with the real-time optimization dynamics of the model and explicitly enhancing memory, emp avoids the distribution mismatch and representation bias that plague static methods. This is crucial for high pruning rates, as static subsets are increasingly inconsistent with the evolving generalization requirements of the model.
@@ 4 Experiment
In the following sections, we validate the effectiveness of our theoretical results and the proposed dataset pruning method through experiments. In Section 4.1, we compare EMP with several other baseline methods on SL, including image classification tasks and a range of natural language tasks. In Section 4.2, we verify the high efficiency of EMP on SSL. In Section , we conduct a series of ablation experiments to validate the effectiveness of our theoretical results. In Section 4.4, we analyze the generalization performance of EMP using a one-dimensional linear interpolation method. It is worth noting that other baseline methods include static pruning: CD [ref30], Herding [ref31], DeepFool [ref32], Last Confidence [ref48], Glister [ref51], and dynamic pruning: InfoBatch [ref4], $\epsilon$-greedy, and UCB [ref5]. It is important to note that EMP is also a dynamic pruning method, so we mainly compare it with other dynamic pruning methods. To eliminate the influence of randomness, each of our experiments was run 5 times, and the average was taken. Our dataset introduction and experimental details are in the Appendix B and C.
@@ 4.1 Performance of EMP in SL
CIFAR. We conducted comparisons on both CIFAR-10 and CIFAR-100. In this work, similar to InfoBatch [ref4], we employed an annealing technique, specifically, using the entire dataset for training towards the end of the training process. More details are shown in Section 4.3. We presented results from other methods in recent years in Table (1). EMP achieved lossless performance on CIFAR-100, even slightly higher than the baseline. Notably, EMP significantly surpassed previous static pruning methods and led other dynamic pruning methods at high pruning rates. Specifically, on the CIFAR-100 dataset, EMP exceeded other methods by at least 1.3% under any pruning rate, and even by at least 2.1% at a 70% pruning rate.
@@
CD([ref30]) 95.0 94.3 90.8 74.2 72.3 70.3
Herding([ref31]) 92.2 88.0 80.1 73.1 71.8 69.6
K-Center([ref47]) 94.7 93.9 90.9 74.1 72.2 70.2
Least Confidence([ref48]) 95.0 94.5 90.3 74.2 72.3 69.8
Margin([ref48]) 94.9 94.3 90.9 74.0 72.2 70.2
GraNd-4([ref49]) 95.3 94.6 91.2 74.6 71.4 68.8
DeepFool([ref32]) 95.1 94.1 90.0 74.2 73.2 69.8
Craig([ref50]) 94.8 93.3 88.4 74.4 71.9 69.7
Glister([ref51]) 95.2 94.0 90.9 74.6 73.2 70.4
EL2N-2([ref3]) 94.4 93.2 89.8 74.1 71.0 68.5
EL2N-20([ref3]) 95.3 95.1 91.9 77.2 72.1 -
DP([ref33]) 94.9 93.8 90.8 77.2 73.1 -
@@
$\epsilon$-greedy([ref5]) 95.2 94.9 94.1 76.4 74.8 75.1
UCB([ref5]) 95.3 94.7 93.9 77.3 75.3 74.8
InfoBatch([ref4]) 95.6 95.1 94.7 78.2 78.1 76.5
[…]
@@
ImageNet-1K. To explore the performance of EMP on a large-scale dataset, we trained ResNet50 on ImageNet-1K, with results as shown in Table 2. Notably, the UCB and Greedy methods demonstrated performance similar to that of dynamic random pruning, and at high pruning rates, dynamic random pruning surpassed them. We believe that on ImageNet-1K, due to some noisy samples [ref34], they are difficult to fit correctly, hence LFL tends to select these noisy samples, leading to suboptimal performance of LFL. As expected, on the challenging ImageNet-1K dataset with high fitting difficulty, EMP showed efficient performance both at low and high pruning rates.
@@
[…]
GLUE. In addition to image classification tasks, natural language understanding tasks are also within the scope of our investigation. Specifically, we used the BERT-base pre-trained model [ref24] for fine-tuning on the GLUE benchmark [ref35]. The experimental results are reported in Table (3), and it is evident that EMP leads in most cases, even outperforming the case without pruning. Notably, at pruning rates of 50% and 70%, for datasets with a larger number of samples such as MNLI (393K) and QQP (363K), EMP demonstrates good performance. Specifically, at a 70% pruning rate, EMP leads other methods by at least 1% on the MNLI dataset. On the dataset QNLI with a moderate number of samples (108K), EMP also shows superior performance, exceeding other methods by 0.6%-1.9% at a 70% pruning rate. This proves that our method is also effective on natural language tasks.
@@ 4.2 Performance of EMP in SSL
Recently, Large Language Models (LLMs) have demonstrated significant performance [ref25, ref23], pretraining on vast datasets to obtain models with general knowledge. Visually, various visual pretrained models [ref36, ref37] are also gaining popularity. To explore the performance of EMP in model pretraining, we experimentally verified it on CL using the SimCLR framework [ref15]. In this work, we specifically pruned the data during the pretraining phase to investigate the practical effectiveness of EMP.
@@
$\epsilon$-greedy([ref5]) 82.01 79.98 76.91 53.11 49.89 42.72
UCB([ref5]) 81.37 79.24 71.85 52.50 49.34 42.15
InfoBatch([ref5]) 83.09 81.39 79.05 55.21 50.56 46.01
[…]
@@ 4.3 Ablation Experiment
It can be seen that the degree of annealing increases with the increase of $\alpha$, but this also brings corresponding additional overhead. We believe that when the model has acquired sufficient knowledge in the non-annealing phase, the positive effects brought by the annealing phase will be less significant. In other words, the annealing phase not only exists as a technique to improve accuracy but also serves as an effective measure of the validity of the data pruning algorithm. Similarly, InfoBatch [ref4] also uses this technique, and we mainly compare it with this method. Our results are reported in Figure (8). It is not difficult to see that at lower pruning rates, both methods are robust to annealing, and when the pruning rate increases, the shortcomings of LFL become apparent. Although EMP’s accuracy drops without the annealing phase, InfoBatch’s drop is more dramatic. Therefore, we can conclude that under any circumstances, EMP is robust to the annealing technique, which means that EMP enables the model to remember more samples and gain more knowledge during the early to mid-training phase.
@@
Label Noise. Label noise is a common challenge in real-world applications. Enhancing the robustness of algorithms to label noise during the data pruning process is a key issue [ref17]. In this section, we investigate the robustness of EMP to label noise by conducting comparative experiments on CIFAR10 and CIFAR100 with synthetic label noise. Specifically, we introduce label noise [ref38] into the two datasets by randomly replacing the labels of a certain percentage of training data with all possible labels, which mimics the real-life scenario where researchers may misjudge images [ref39]. Our results are reported in Figure (9). It can be observed that EMP still maintains high performance even with the addition of 20% noise and leads other methods. Notably, the Greedy and UCB methods are not robust to noise, and are even outperformed by random dynamic pruning. We believe that these noisy data are difficult to fit during training, hence their scores remain high, leading to the retention of noisy data and resulting in low model performance. Overall, EMP maintains high performance in the presence of noise, indicating that EMP indeed retains key or general samples, thus demonstrating robustness to noise.
@@ 4.4 Generalization Analysis
To further investigate the effectiveness of EMP, we analyze its generalization capability. We use one-dimensional linear interpolation to examine how EMP affects the loss landscape while enhancing the model’s memory ability. Previous work [ref41, ref42] suggests that better model generalization is associated with flat minima in the loss landscape. Following the method proposed in [ref40], we inspect the loss landscape. Specifically, we assess the performance of models with parameters $(1-\epsilon)\theta_{\text{init}}+\epsilon\theta_{T}$, where $\theta_{\text{init}}$ is the initial model, and $\theta_{T}$ is the converged model after optimization with different dataset pruning algorithms.
@@
4.5 Class Imbalance
To investigate how EMP addresses class imbalance, we conducted an experimental study. The ten classes of CIFAR-10, namely ’airplane’, ’automobile’, ’bird’, ’cat’, ’deer;, ’dog’, ’frog’, ’horse’, ’ship’, and ’truck’, are denoted as C1 to C10 respectively in the following sections of the work. As shown in Figure  12 and 13, the data subset selected by EMP exhibits significant class imbalance?one class accounts for nearly 20% of the samples in CIFAR-10. Vysogorets et al.[vysogoretsdrop] point out that the key to effective pruning lies in an appropriate difficulty-based class ratio rather than strict class balance.
Therefore, we designed a new experiment: using EMP?s sample scoring function while enforcing class balance. The corresponding performance results are reported in Table 6. It can be observed that the performance of the EMP method significantly deteriorates after class balancing, approaching results comparable to random pruning. This outcome aligns with findings from other methods, indicating that EMP does not enhance performance by resolving class imbalance but rather by identifying samples that genuinely benefit model performance and generalization.
In addition, we designed a new experiment by redistributing class proportions according to the ratios in Table 7 and trained models under this configuration. The results are shown in Figure  14. We observe that under high pruning rates, all methods (including EMP) fail to maintain accuracy, as certain sample features cannot be effectively learned. However, we find that EMP outperforms other methods. We argue that in LFL methods (represented by InfoBatch), the model?s higher uncertainty when predicting minority classes leads to more frequent selection of these classes under high pruning rates. Yet, due to missing critical features, these samples often become hard-to-fit. We report this phenomenon in Figure 15. In contrast, EMP mitigates this limitation via its memory term by selecting more representative samples for training. Through this mechanism, EMP effectively alleviates class imbalance issues.
Figure 12: Under supervised learning, we evaluate the percentage of retained samples per class for a ResNet-18 model on the CIFAR-10 dataset at a pruning ratio of 70%
Figure 13: Under self-supervised learning, we evaluate the percentage of retained samples per class for a ResNet-18 model on the CIFAR-10 dataset at a pruning ratio of 70%
Figure 14: Comparative results of different methods under varying pruning rates using the ResNet10 model and a class-imbalanced CIFAR10 dataset with the class proportions specified in Tab. 7
Figure 15: In the InfoBatch method, we evaluated the per-class retention percentage of samples for the ResNet-18 model on the class-imbalanced CIFAR-10 dataset at a pruning rate of 70%
Table 6: Comparative results of EMPbalance (with enforced equal class proportions) versus other methods on ResNet18 at a 50% pruning rate
EMPbalance EMP DP InfoBatch Randomdynamic
CIFAR10 94.88 95.23 93.8 95.1 94.5
CIFAR100 72.21 79.44 73.1 78.1 72.1
Table 7: The subsampling rates by class for the imbalanced CIFAR10 dataset
Class Indices Subsample Rate
c1, c2 25%
c3, c4, c5 50%
c6, c7 75%
c8, c9, c10 100%